# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ 2.1 Our attack algorithm
To address the challenge of non-unique solutions in gradient matching, we introduce a regularization term designed to align intermediate network features with those reconstructed via a tensor-based method. This regularization, denoted as $\mathcal{R}_{\text{feature}}(x)$, employs squared cosine similarity to mitigate the issue of sign ambiguity inherent in tensor-based reconstruction:
$\mathcal{R}_{\rm feature}({x})=1-\left(\frac{\langle f({x},\phi),\hat{z}\rangle}{\left\|f({x},\phi)\right\|\left\|\hat{z}\right\|}\right)^{2},$ (2)
where $f(x,\phi)$ are the intermediate features of input $x$ and $\hat{z}$ are the reconstructed features from the tensor-based method. This approach aims to refine gradient matching by ensuring the generated input $x$ yields intermediate features similar to $\hat{z}$, providing a more accurate and specific solution space.
@@
3 Defense Methods
To combat privacy attacks of machine learning, some defenses are proposed to prevent data leakage. Specifically in the setting of data reconstruction attack in federated learning, (Zhu et al., 2019; Geiping et al., 2020; Huang et al., 2021) evaluated the privacy of different defending strategies empirically. These works shed light on choosing appropriate defenses against reconstruction attacks but can be improved in the following sense: (i) the attacks they used to evaluate defenses lack strength. However, the definition of how robust a defense strategy is should be measured on the strongest attack within the threat model; and (ii) previous evaluations are purely empirical without theoretical support. In contrast, our evaluation of defenses is based on a stronger attack we proposed. Moreover, we also provide theoretical analysis from the perspective of feature reconstruction.
Our attack is composed of gradient matching and feature matching. Defenses bring both optimization and statistical barriers to the gradient matching component, where the effect is difficult to analyze theoretically. The feature-matching component, on the other hand, can be analyzed quantitatively. The quality of feature matching is not only essential to and correlates strongly with the overall performance of data reconstruction, but the intermediate feature, as well as original data, is also protective information and its leakage should be prevented as well. To this end, we will theoretically analyze the defending effect on feature reconstruction in the following part of this section.
With our stronger attack, the feature reconstruction is equivalent to recovering the input of a two-layer neural network. Thus, our theoretical analysis can be based on the results of reconstruction error bounds on two-layer networks in (Wang et al., 2023) but varies among different defenses. On one hand, our theoretical findings overturned some prior empirical observations on the effectiveness of several defense methods. On the other hand, it is generally more convincing to analyze the effect of defenses under a stronger and more robust attacking method. Combined with the experiments in Section 4, our analysis gives a more complete understanding of various defending methods.
In the following analysis, our setting conforms to that of Section 2.1 with a main focus on feature reconstruction of fully connected layers.
3.1 No defense
Without any defense, we conduct tensor decomposition to $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j})$ as stated in Section 2.1. Then by the following Stein’s lemma (Stein, 1981; Mamis, 2022):
Lemma 3.1 (Stein’s Lemma).
$\mathbb{E}[g(X)H_{p}(X)]=\mathbb{E}[g^{(p)}(X)],$ (4)
we have
$\sum_{j=1}^{m}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j})\approx$ $m\operatorname{\mathbb{E}}_{{\bm{w}}\sim\mathcal{N}(0,I)}[g({\bm{w}})H_{p}({\bm{w}})]$
$=m\operatorname{\mathbb{E}}_{\bm{w}}[\nabla^{p}_{\bm{w}}g({\bm{w}})]=$ $\sum_{i=1}^{B}r_{i}\mathbb{E}[\sigma^{(p)}({\bm{w}}^{\top}{\bm{x}}_{i}){\bm{x}}_{i}^{\otimes p}].$ (5)
Thus, the result of tensor decomposition is $\{x_{1},\dots,x_{B}\}$ and the error bound is $\tilde{O}(BK\sqrt{\frac{d}{m}})$ with high probability, where $B$ is the batch size of data. Details of the analysis is in Appendix B.
3.2 Local aggregation
In a realistic setting of federated learning, the observed local update may not be a single gradient. Instead, it can be the result of multiple steps of gradient descent. Specifically, under the setting of federated averaging (McMahan et al., 2017), a client with $N$ images trains $E$ epochs for each update, where a single epoch contains $N/B$ gradient descent steps with batch size $B$.
Since the observation is more complicated than a gradient, the difficulty of reconstructing features increases. However, the tensor-based attack can still work under this setting. For an update $\tilde{g}$ with multiple steps, we can simply replace the gradient $g$ in the previous attack with $\tilde{g}$ and conduct tensor decomposition as normal. Specifically, we will discuss two cases: using the same batch to train multiple steps and using different batches for every gradient step.
Multiple steps with the same batch.
In this setting, $B=N$ and $E>1$. We first consider the case $E=2$. According to gradient descent, we have $a_{kj}^{(t+1)}=a_{kj}^{(t)}-\frac{\eta}{m}\sum_{i=1}^{B}(p_{ki}^{(t)}-y_{ki}^{(t)})\sigma({\bm{w}}_{j}^{(t)}\cdot{\bm{x}}_{i})$ for $t=0,1$, where $p_{ki}^{(t)},y_{ki}^{(t)},a_{kj}^{(t)},{\bm{w}}_{j}^{(t)}$ are the values of corresponding parameters after $t$ gradient steps. Since the information from clients are the values of $a_{j}^{(2)}$, we have from the local update
$\begin{split}a_{kj}^{(2)}-a_{kj}^{(0)}=-\frac{\eta}{m}\sum_{i=1}^{B}(p_{ki}^{(0)}-y_{ki}^{(0)})\sigma({\bm{w}}_{j}^{(0)}\cdot{\bm{x}}_{i})-\frac{\eta}{m}\sum_{i=1}^{B}(p_{ki}^{(1)}-y_{ki}^{(1)})\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i}).\end{split}$ (6)
Under this setting, we can compute $\sum_{j=1}^{m}\sum_{k=1}^{K}c_{k}(a_{kj}^{(2)}-a_{kj}^{(0)})H_{p}({\bm{w}}_{j})$ and conduct rank-$B$ tensor decomposition, which is similar to the original problem. For the attack, we can prove that the reconstruction error is bounded by $\tilde{O}(BK\sqrt{\frac{d}{m}})$ with high probability.
Different batches.
In this setting, $E=1$ but $N>B$. We first consider the case with only 2 batches, i.e. $B<N\leq 2B$. Similar to Eq.(6), the local update
$\begin{split}a_{kj}^{(2)}-a_{kj}^{(0)}=&-\eta\sum_{i=1}^{B}(p_{ki}^{(0)}-y_{ki}^{(0)})\sigma({\bm{w}}_{j}^{\top}{\bm{x}}_{i})-\eta\sum_{i=B+1}^{N}(p_{ki}^{(1)}-y_{ki}^{(1)})\sigma({\bm{w}}_{j}^{(1)}\cdot{\bm{x}}_{i}).\end{split}$ (7)
The only difference from the case above is that the data used to train $\Theta^{(1)}$ is different from the one in training $\Theta^{(0)}$. Thus, performing our attacking method with $N$ outputs instead of $B$ can reconstruct the original data. We can prove that the reconstruction error is bounded by $\tilde{O}(NK\sqrt{\frac{d}{m}})$ with high probability.
Conclusion.
Note that for both cases, if there are more than 2 gradient steps, we can perform the same tensor decomposition method. Thus, the tensor-based attack can recover features against local aggregation attacks. The proofs of the error bounds above are in Appendix C.1.
3.3 Differential privacy stochastic gradient descent
Differential privacy (Dwork, 2006) algorithms ensure the privacy of data by adding noise to gradient descent steps (Shokri and Shmatikov, 2015; Abadi et al., 2016; Song et al., 2013). In this way, it hides the exact local gradient from the global model. In differential private federated learning (Geyer et al., 2017; Wei et al., 2020b), each client clips the gradient and introduces a random Gaussian noise before updates to the cluster. In the setting of two-layer networks, the update of $a_{j}$ is
$\tilde{g}_{j}=g_{j}/\max\left\{1,\frac{\|g\|}{S}\right\}+\epsilon,\text{ where}\epsilon\sim\mathcal{N}(0,\sigma^{2}I).$
Note that under this setting, the information contained in the gradient is disrupted by gradient clipping and random noise. We first analyze the effect of the two strategies separately. For gradient clipping, we conduct our method based on observed clipped gradient $\tilde{g}$. There will be no defensive effect to our attack no matter the clipping threshold and we can prove that the reconstruction error bound $\tilde{O}(BK\sqrt{\frac{d}{m}})$ is the same as that with no defense (even for the hidden terms). For noisy gradient $g+\epsilon$, we have the same order of error bound if the noise level $\sigma\leq O(\frac{B}{m})$ so the defense is not effective.
In the defense of differential privacy gradient descent, we combine the two strategies and conduct a tensor method based on $\tilde{g}+\epsilon$. The defense is more effective and has an error bound $O(\max\{1,\frac{\|g\|}{S}\})$ times to the bound of gradient noise defense only. Gradient clipping enhances the defensive performance of the gradient noise though itself has no effect in defending the attack. However, our attack method is still enough to recover the feature since we can modify the initial weights to control $\|g\|$.
Conclusion.
The defense of gradient clipping has no effect while gradient noise and differential privacy gradient descent is successful only when the noise level is relatively large. However, the noise variance $\sigma^{2}$ in the application is usually small since a large random noise will cause bad performance in federated learning (Wei et al., 2020b). We also show some results in Section 4. See detailed proofs of the error bounds in Appendix C.2.
3.4 Secure aggregation
In federated learning, local devices update gradients to the server individually. Secure aggregation (Bonawitz et al., 2016, 2017) is a method that clients $\{u_{1},\dots,u_{L}\}$ can aggregate their gradients prior to the cluster and the cluster can only access to aggregated gradient: $\tilde{G}=\frac{1}{B}\sum_{i=1}^{L}g_{l}B_{l},$ where $g_{l}$ and $B_{l}$ is the gradient update and batch size for user $u_{l}$ respectively and $B=\sum_{i=1}^{L}B_{l}$. In this way, the global server is blocked from knowing each individual gradient and then prevents privacy leakage. Even though the server can reconstruct the training data, it cannot identify which user has the data. If we only reconstruct data without identifying its source, the aggregated gradient is equivalent to a gradient step with large batch size $B$ and we can conduct our tensor-based attack on the aggregated gradient.
Conclusion.
Tensor method will still work when the batch size is not too large. Otherwise, say $B>d$, feature recovery cannot work since the tensor decomposition is not unique.
3.5 Dropout
Another method to improve the defending effect is dropout (Hinton et al., 2012; Srivastava et al., 2014), which is a mechanism designed to prevent overfitting. It randomly drops nodes in a network with a certain probability $1-p$. In this way, some of the entries in the gradient randomly turn to zero. Since dropout introduces randomness to the gradient, it can be used to defend against privacy attacks in federated learning. The effect of dropout against gradient matching is empirically discussed in (Zhu et al., 2019; Huang et al., 2021). A two-layer fully connected network with a dropout layer can be written as
$f({\bm{x}};\Theta)=\sum_{j=1}^{m}s_{j}a_{j}\sigma({\bm{w}}_{j}\cdot{\bm{x}})=\sum_{j:s_{j}=1}a_{j}\sigma({\bm{w}}_{j}\cdot{\bm{x}}),$
where $s_{j}\sim\mathrm{Bernoulli}(p)$. Thus, the model has an effective width equal to the number of nodes $m^{\prime}$ that has not been dropped. We can compute the sum $\frac{1}{m^{\prime}}\sum_{j:s_{j}=1}g_{j}H_{p}({\bm{w}}_{j})$ and conduct tensor decomposition. Since the effective width $m^{\prime}\approx pm$, the defense is not effective when $p$ is not too small.
3.6 Gradient pruning
Gradient pruning (Sun et al., 2017; Ye et al., 2020) is a method that accelerates the computation in training by setting the small entries in gradient to zero. Thus, the gradient will be sparse, which is similar to dropout. However, the key difference is that the remaining entries of the gradient in dropout are chosen randomly yet gradient pruning drops small entries making the defending more effective (when a similar level of utility is lost).
If we denote the set of entries that is not pruned by $S$ and conduct the tensor-based attack, we have
$\hat{{\bm{T}}}_{p}=\frac{1}{m^{\prime}}\sum_{j\in S}g_{j}H_{p}({\bm{w}}_{j}),$
where $g_{j}=\nabla_{a_{j}}L$ and $m^{\prime}=|S|$. This formula has a similar form to the case in dropout but the key difference is that the remaining ${\bm{w}}_{j}$’s are more likely to make $g_{j}$ larger so they do not follow the normal distribution. In this case, the concentration property $\sum_{j\in S}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j})\approx m^{\prime}\operatorname{\mathbb{E}}_{{\bm{w}}\sim\mathcal{N}(0,I)}[g({\bm{w}})H_{p}({\bm{w}})]$ may not hold. Moreover, Lemma 3.1 cannot be applied with ${\bm{w}}_{j}$’s are not normal random vectors so $\operatorname{\mathbb{E}}_{{\bm{w}}\sim\mathcal{N}(0,I)}[g({\bm{w}})H_{p}({\bm{w}})]\neq\operatorname{\mathbb{E}}_{\bm{w}}[\nabla^{p}_{\bm{w}}g({\bm{w}})]$. Thus, the $\hat{{\bm{T}}}_{p}$ is not a good estimation of the real tensor product, indicating gradient pruning defense is effective against feature reconstruction.
@@ 4 Experiments
Batch Size 16 32
Method Geiping Jeon Ours Geiping Jeon Ours
LPIPS normal-$\downarrow$ 0.41 (0.09) 0.17 (0.12) 0.14 (0.11) 0.45 (0.11) 0.24 (0.13) 0.15 (0.11)
Table 1: Comparison of our methods with other state-of-the-art methods, Geiping, and Jeon refer to (Geiping et al., 2020) and (Jeon et al., 2021) respectively. We highlight the best performances in bold.
@@ 4.1 Experimental setup
We evaluate defense methods on CIFAR-10 dataset (Krizhevsky et al., 2009) with a ResNet-18 (trained for one epoch), which is the default backbone model for federated learning. For the deep image prior model, we use a U-Net backbone similar to an unmasked PixelCNN++ (Salimans et al., 2017; Ronneberger et al., 2015) and replace the original weight normalization (Salimans and Kingma, 2016) with group normalization (Wu and He, 2018). We use self-attention at the $16\times 16$ feature map resolution (Wang et al., 2018). The details are in the Appendix D. For GradPrune: gradient pruning sets gradients of small magnitudes to zero. We vary the pruning ratio $p$ in $\{0.3,0.5,0.7,0.9,0.99\}$. For GradClipping: gradient clipping set gradient to have a bounded norm, which is ensured by applying a clipping operation that shrinks individual model gradients when their norm exceeds a given threshold. We set this threshold as $\{2,4,8\}$. We also adopt defense by adding noise to gradient, the noise is generated from random Gaussian with different standard deviations $\{0.001,0.01,0.05,0.1\}$. Local gradient aggregation aggregates local gradient by performing a few steps of gradient descent to update the model. We locally aggregate $3$ or $5$ steps.
@@
We use a subset of 50 CIFAR-10 images to evaluate the attack performance. To achieve a strong attack, we add reconstructed features (after linear projections to match the dimension) into each residual block of the U-Net. We also assume the batch norm statistics are available following (Geiping et al., 2020). We observe the difficulty of optimization when using cosine-similarity as the gradient inversion loss under our setting, and therefore address this issue by reweighting the gradients by their $\ell_{0}$ norm. See details in the Appendix DGeiping et al. (2020); Zhu et al. (2019) have shown that small batch size is vital for the success of the attack. We intentionally evaluate the attack with two small batch sizes $4$ and $8$ and two small but realistic batch sizes, 16 and 32.
@@ 4.2 Experiment results
Results of the proposed attack method.
Our method can be easily added to previous methods (Geiping et al., 2020; Zhu et al., 2019). In Table  1, we compare the state-of-the-art methods with the proposed attacking method, our method outperforms previous methods. We visualize reconstructed images in Figure 2. Without any defense, both (Geiping et al., 2020) and our attack method can recover images well from the gradient, and our method produces higher-quality images. Increasing the batch size makes all methods more difficult to recover well. With defenses, our method shows better robustness – the recovered images are visually similar to the original image. This result shows that our attack is a stronger attack under these defense methods. Figure 2(b) suggests that the feature matching loss is the main reason behind such robustness, showing that the analysis of feature reconstruction is reasonable. See Table 3 and Table 5 in Appendix E for summary of reconstruction results.
@@
Batchsize w/o defense p=0.3 p=0.5 p=0.7 bound=2 bound=4 bound=8 0.001 0.01 0.1 step=3 step=5
@@
Table 2: Feature reconstruct error under different defenses. When using gradient pruning defense and gradient noise defense with large $\sigma$, the feature reconstruction quality degrades significantly.
None GradNoise ($\sigma$) GradPrune ($p$) GradClipping (bound) Local Aggregation (step)
Parameter - 0.001 0.01 0.05 0.1 0.3 0.5 0.7 0.9 0.99 2 4 8 3 5
@@
RMSE normal-$\downarrow$ 0.15(0.04) 0.17 (0.05) 0.22 (0.07) 0.31 (0.06) 0.28 (0.08) 0.16 (0.08) 0.17 (0.07) 0.20 (0.09) 0.26 (0.09) 0.27 (0.10) 0.17 (0.08) 0.17 (0.08) 0.19 (0.08) 0.23 (0.09) 0.25 (0.10)
PSNR normal-$\uparrow$ 23.69 (5.69) 23.31 (3.04) 21.32 (3.52) 14.58 (1.59) 13.11 (4.26) 22.63 (6.15) 23.33 (6.03) 19.40 (6.02) 16.39 (6.56) 14.72 (6.45) 23.26 (6.10) 23.40 (6.00) 23.01 (6.28) 18.65 (6.75) 18.77 (7.33)
@@
RMSE normal-$\downarrow$ 0.15 (0.03) 0.19 (0.07) 0.24 (0.08) 0.29 (0.03) 0.27 (0.07) 0.16 (0.04) 0.16 (0.03) 0.21 (0.04) 0.27 (0.07) 0.27 (0.08) 0.16 (0.03) 0.16 (0.04) 0.16 (0.04) 0.28 (0.08) 0.26 (0.08)
PSNR normal-$\uparrow$ 23.59 (5.39) 20.12 (3.58) 16.41 (4.64) 11.27 (0.33) 12.28 (4.08) 23.16 (5.48) 24.08 (5.38) 18.73 (5.63) 13.90 (3.77) 12.43 (2.73) 22.88 (5.31) 23.93 (5.55) 24.04 (5.58) 13.84 (5.38) 14.21 (5.41)
@@
RMSE normal-$\downarrow$ 0.15 (0.04) 0.19 (0.05) 0.29 (0.05) 0.30 (0.03) 0.30 (0.06) 0.15 (0.03) 0.16 (0.04) 0.20 (0.05) 0.29 (0.05) 0.29 (0.05) 0.16 (0.03) 0.16 (0.04) 0.16 (0.04) 0.29 (0.04) 0.30 (0.04)
PSNR normal-$\uparrow$ 23.37 (5.42) 20.31 (4.13) 14.56 (0.88) 11.27 (0.89) 11.27 (1.83) 23.74 (5.21) 23.32 (5.13) 18.25 (5.22) 11.30 (2.10) 11.13 (2.15) 23.36 (5.14) 23.59 (5.41) 23.47 (5.23) 11.25 (2.56) 10.82 (2.20)
(a) Feature reconstruction error vs final training loss.
(b) PSNR of our method against different defenses.
Figure 3: The results of defenses against our attack. (a) The final training loss measures the level of interference with the training of defenses. The relation between feature error and interference reflects the effect of defenses. (b) PSNR of reconstructed data using the method with and without feature matching. The difference indicates that feature matching improves robustness against defenses.
Comparison of defense methods in feature recovery.
For feature recovery, we compared the reconstruction error under different defending strategies, and the results are shown in Table 2. We also use the final training loss of the federated learning model to measure the level of interference from defenses toward the training (details see Table 6 in Appendix E). For an effective defense, it should increase the feature reconstruction error while disturbing little to the training. We show the relation between reconstruction error and final training loss with different defenses in Figure 2(a). Gradient pruning has a large reconstruction error without increasing training loss too much so it has the best defending effect under this criteria. Adding noise also prevents feature leakage when noise is large enough but it will affect training. Gradient clipping is barely effective for defending feature recovery while dropout has a mild effect but it is harmful to training. The feature reconstruction errors are consistent with the theoretical analysis in Section 3.
Comparison of defense methods in inputs reconstruction.
According to 3, data reconstructed by our method with defense gradient pruning and gradient noise has the smallest average PSNR, indicating that these two defenses are effective against our attack. However, according to Figure 2(a), the final training loss of gradient pruning is smaller than that of gradient noise even when the pruning parameter $p$ is very large. Therefore, gradient pruning is the best defense that prevents data leakage and maintains training utility at the same time. It is also shown in Figure 2 that gradient pruning with $p=0.9$ has the best defensive effect. Gradient noise with $\sigma=0.1$ also works but the final training loss is large, showing its interference towards training.
We also compared between our attack method and the method without feature matching against different defenses in Figure 2(b) to show the effect of feature matching. Here we choose the batch size $B=8$ and defenses gradient noise with $\sigma=0.001$, gradient pruning with $p=0.9$, and gradient clipping with threshold 4. It is shown that feature matching improves the robustness of our attack against gradient noise and gradient clipping defenses. This can be explained by our analysis in Section 3 that feature recovery works badly against gradient pruning and feature matching regularization has no effect with inaccurate features. See the quantified effect of defenses with and without feature matching in Table 3 and Table 5 in Appendix E.

### v2

@@ 4.2 Description of the proposed attack algorithm
By treating the intermediate (last but one) layer ${\bm{z}}_{i}$ as the model input, we approximately reconstruct $\hat{z}_{i}$ as a preprocessing step using the tensor-based method in (Wang et al., 2023). However, the reconstructed quality, the ordering of $\hat{z}_{i}$, or whether it is unique is unclear. To address such challenges, we introduce a gradient-matching regularization term to align intermediate network features with reconstructed feature $\hat{z}_{i}$. This regularization, denoted as $\mathcal{R}_{\text{feature}}(x)$, employs squared cosine similarity to mitigate the issue of sign ambiguity inherent in tensor-based reconstruction: $\mathcal{R}_{\rm feature}({x})=1-\left(\frac{\langle f({x},\phi),\hat{z}\rangle}{\left\|f({x},\phi)\right\|\left\|\hat{z}\right\|}\right)^{2},$ where $f(x,\phi)$ are the intermediate features of input $x$ and $\hat{z}$ are the reconstructed features from the tensor-based method. This approach aims to refine gradient matching by ensuring the generated input $x$ yields intermediate features similar to $\hat{z}$, providing a more accurate and specific solution space. We notice the reconstruction quality of individual samples drops drastically when they fall into the same class. However, the reconstruction of their spanned space is still accurate. We further refine the regularize as $\mathcal{R}_{\rm feature}({x})=\|P_{\text{span}(\hat{z}_{i})}^{\perp}f(x,\phi)\|^{2}$ when the sample size is large.
@@ 5.1 Experimental setup
We evaluate defense methods on CIFAR-10 dataset  (Krizhevsky et al., 2009) with a ResNet-18 (trained for one epoch), which is the default backbone model for federated learning. The details are in the Appendix F. For GradPrune: gradient pruning sets gradients of small magnitudes to zero. We vary the pruning ratio $p$ in $\{0.3,0.5,0.7,0.9,0.99\}$. For GradClipping: gradient clipping set gradient to have a bounded norm, which is ensured by applying a clipping operation that shrinks individual model gradients when their norm exceeds a given threshold. We set this threshold as $\{2,4,8\}$. We also adopt defense by adding noise to gradient, the noise is generated from random Gaussian with different standard deviations $\{0.001,0.01,0.05,0.1\}$. For GradDropout, we consider randomly dropout the entries of the gradient with probability in $\{0.3,0.5,0.7,0.9\}$. Local gradient aggregation aggregates local gradient by performing a few steps of gradient descent to update the model. We locally aggregate $3$ or $5$ steps.
@@
We use a subset of 50 CIFAR-10 images to evaluate the attack performance. For all attack algorithms, we assume the batch norm statistics are available following (Geiping et al., 2020). We observe the difficulty of optimization when using cosine-similarity as the gradient inversion loss for the proposed attack algorithm and therefore address this issue by reweighting the gradients by their $\ell_{0}$ norm. See details in the Appendix FGeiping et al., 2020; Zhu et al., 2019 have shown that small batch size is vital for the success of the attack. We intentionally evaluate the attack with two small batch sizes $4$ and $8$ and two small but realistic batch sizes, 16 and 32.
@@ 5.2 Experiment results
Batchsize w/o defense $p=0.3$ $p=0.5$ $p=0.7$ $C=2$ $C=4$ $C=8$ 0.001 0.01 0.1 step=3 step=5
@@
Table 2: Feature reconstruction error under different defenses. When using gradient pruning defense and gradient noise defense with large $\sigma$, the feature reconstruction quality degrades significantly.
Verification of theoretical analysis.
In Section 3, we analyze the performance of different defenses on 2-layer networks. In the experiments, our attack method involves intermediate feature matching, which requires reconstructing feature from a 2-layer network. This setting aligns with our theoretical analysis. Therefore, we can verify the theory by checking the feature reconstruction error in our attack. The results is shown in Table 2, where the reconstruction error increases when the batch size increases for all setting. This matches the theoretical bounds. Moreover, gradient pruning and gradient noise with large scaling has significantly large reconstruction error, which also aligns with our theoretical evaluation.
Batch Size 16 32
Method Geiping Jeon Ours Geiping Jeon Ours
LPIPS $\downarrow$ 0.41 (0.09) 0.17 (0.12) 0.14 (0.11) 0.45 (0.11) 0.24 (0.13) 0.15 (0.11)
Table 3: Comparison of our methods with other methods, Geiping, and Jeon refer to (Geiping et al., 2020) and (Jeon et al., 2021) respectively. We highlight the best performances in bold.
Our proposed attack.
According to the framework of evaluation, we need to consider strong attacks against various defenses. A key motivation of our proposed attack is its good performance among different defense, which we will verify in the following results. Our method is based on gradient inversion and can be easily added to previous methods (Geiping et al., 2020; Zhu et al., 2019). In Table  3, we compare the state-of-the-art gradient inversion methods with the proposed attacking method, our method outperforms previous methods. We visualize reconstructed images in Figure 2. Without any defense, both (Geiping et al., 2020) and our attack method can recover images well from the gradient, and our method produces higher-quality images. Increasing the batch size makes all methods more difficult to recover well. With defenses, our method shows better robustness – the recovered images are visually more similar to the original image. See Table 7 and Table 8 in Appendix I for summary of reconstruction results.
GradClip ($C$) GradDrop ($p$) GradNoise ($\sigma_{0}$) GradPrune ($p$)
Parameter 2 4 8 0.3 0.5 0.7 0.9 0.001 0.01 0.1 0.3 0.5 0.7 0.9 0.99
$B=2$ Ours 0.17 0.17 0.19 0.16 0.16 0.18 0.18 0.17 0.22 0.28 0.16 0.17 0.20 0.26 0.27
GradientInversion 0.19 0.23 0.25 0.19 0.20 0.20 0.21 0.30 0.32 0.35 0.19 0.20 0.24 0.28 0.27
Robbing The Fed 0.23 0.23 0.23 0.28 0.32 0.32 0.30 0.26 0.25 0.27 0.22 0.25 0.32 0.37 0.46
CPA 0.21 0.22 0.23 0.23 0.22 0.22 0.24 0.22 0.23 0.28 0.21 0.23 0.29 0.28 0.32
$B=4$ Ours 0.16 0.16 0.16 0.16 0.16 0.19 0.19 0.19 0.24 0.27 0.16 0.16 0.21 0.27 0.27
GradientInversion 0.21 0.22 0.23 0.21 0.21 0.21 0.22 0.31 0.31 0.31 0.19 0.18 0.23 0.29 0.28
Robbing The Fed 0.18 0.18 0.18 0.27 0.31 0.32 0.31 0.19 0.19 0.22 0.20 0.25 0.32 0.39 0.42
CPA 0.21 0.21 0.20 0.21 0.23 0.22 0.24 0.23 0.23 0.25 0.18 0.21 0.24 0.29 0.31
$B=8$ Ours 0.16 0.16 0.16 0.16 0.17 0.19 0.19 0.19 0.29 0.30 0.15 0.16 0.20 0.29 0.29
GradientInversion 0.22 0.21 0.21 0.22 0.22 0.22 0.23 0.30 0.31 0.31 0.20 0.21 0.25 0.30 0.30
Robbing The Fed 0.16 0.16 0.16 0.29 0.31 0.31 0.31 0.16 0.16 0.25 0.22 0.30 0.30 0.36 0.43
CPA 0.21 0.21 0.22 0.22 0.22 0.23 0.25 0.20 0.22 0.24 0.19 0.22 0.26 0.32 0.33
Table 4: We run attacks under different defenses with various parameters and batch sizes. For each setting, we compared the attacks and select the one with lowest RMSE (the highlighted numbers). The best attack for each setting represents the degree of data leakage of that specific defense. Compared with the utility loss of this defense, we can systematically evaluate defenses.
GradClip ($C$) GradDrop ($p$) GradNoise ($\sigma_{0}$) GradPrune ($p$)
Parameter 2 4 8 0.3 0.5 0.7 0.9 0.001 0.01 0.1 0.3 0.5 0.7 0.9 0.99
@@
Final training loss $\downarrow$ 0.390 0.481 0.377 0.466 0.363 0.564 0.864 0.445 0.769 1.889 0.585 0.430 0.621 0.763 1.020
@@
Final training loss $\downarrow$ 0.408 0.353 0.329 0.552 0.316 0.275 0.460 0.527 0.551 1.764 0.253 0.389 0.561 0.505 0.935
@@
Final training loss $\downarrow$ 0.377 0.201 0.357 0.259 0.178 0.231 0.461 0.365 0.343 1.540 0.256 0.161 0.257 0.344 0.725
Table 5: The final training loss of the model with defenses, which measures the utility loss from defenses. Gradient noise produces much larger interference than gradient pruning.
Systematic evaluation of defense.
In our systematic evaluation of different defenses, we measure their strength with the strongest attack: $\mathcal{S}_{D}=\max_{A\in\mathcal{A}}d(S,A(D(G)))$. In order to estimate the maximum reconstruction error, we selected four different attacks to conduct the experiment. The reconstruction RMSE of the four attacks against defenses with various parameters and batch sizes are shown in Table 4. For each setting, we select the attack with the smallest RMSE as the strongest attack. Therefore, among all defenses, gradient pruning with large $p$ and gradient noise with large variance $\sigma_{0}^{2}$ has the best defensive effect.
Figure 3: The relation between the strength of the defense and the utility loss, where we use the best RMSE and the final training loss of the original task respectively. Each dot represents a defense method with a different batch size and strength. For the same level of utility loss, gradient pruning has the best defending effect.
In order to evaluate the defenses, another key criteria is the utility loss of the defense methods. Most of the defenses prevent models from data leakage by perturbing the gradients in training, which may hurt the training task itself. In our experiment, we measure the utility by the final loss of the training task. We show the final training loss for the defenses in Table 5. Gradient pruning, as one of the defenses having strongest effect towards attacks, causes smaller utility loss than gradient noise. This result indicates that gradient pruning is a better defense than gradient noise. In Figure 3, we plot the relation between the reconstruction error and the utility loss, where each dot represents a method under the setting of a specific batch size and parameter of the defense. With the same level of utility loss, gradient pruning has the largest RMSE, indicating it is the best defense to have $\min_{D\in\mathcal{D}_{\mathcal{U}}}\mathcal{S}_{D}$.
Note that there are various of attacks to reconstruction input data and we cannot try all of them on defenses. However, within our framework of evaluation, newly proposed attacks can be added in the experiment and will improve the strength of defenses $\mathcal{S}_{D}$. Therefore, we call for more empirical analysis on top of our results with more attack methods, in order to produce a more accurate evaluation.